Year                                             Cash Flow

Year 0 (Today)                       $2,000

Year 1                                           2,000

Year 2                                                  0

Year 3                                            1,500

Year 4                                            2,500

Year 5                                            4,000

What is the future value (Year 5) of the cash flow stream if the cash flows are invested in an account that pays 10% annually?
Answer #1

Future value of cash flow can be calculated using the basic time value of money function: FV = PV * (1 + r)n

Now, based on this function,

FV = 2000 * (1 + 10%)5 + 2000 * (1 + 10%)4 + 0 * (1 + 10%)3 + 1500 * (1 + 10%)2 + 2500 * (1 + 10%)1 + 4000 * (1 + 10%)0

FV = 3,221.02 + 2,928.20 + 0 + 1815.00 + 2,750 + 4,000

FV = $14,714.22 --> Answer